OpenStudy (anonymous):

Please help... Find all zeros of f(x) = x^3 - x^2 + x -21 and write a linear factorization of f(x). Zeros: ? f(x) = ?

OpenStudy (amriju):

x=3 is a soln...so x-3 is a factor...divide f(x) by (x-3)...u get a quad eqn..solve it...find all roots...hence factorise it...in case the quad eqn returns imaginary roots...u have two factors and 1 zero...else..three factors, 3 zeros

OpenStudy (campbell_st):

in a polynomial f(x) ax^3 + bx^2 + cx + d you need to find the factors of a and d \[a = \pm 1 d = \pm 1, \pm 3, \pm 7.. or .. \pm 21\] let the factors of d = p and factors of a = q then a root of the polynomial will be p/q you just need to find it so (x - 3) is a factor because f(3) = 3^3 -3^2 + 3 - 21 f(3) = 0 This is the factor theorem so if you use polynomial division or synthetic division  you'll find the other factor which will be a quadratic... which will have irrational roots.